### v2

@@ 2 Lorentz-Covariant M2-Brane Matrix Model with RVPD Gauge Symmetry
In this section, we review the Lorentz-covariant M2-brane matrix model with Restricted Volume-Preserving Deformations (RVPD), which we proposed in earlier work[20, 21]. This review will serve as the basis for the BRST gauge-fixing procedure introduced in the next section.
2.1 Notation and Conventions
Throughout this paper we adopt the following notations:
@@ 2.3 Reformulation Using the Nambu Bracket
where $Q_{1},\ Q_{2}$ are arbitrary charges. The Poisson bracket property generalizes to the Nambu bracket, with the Jacobi identity replaced by the Fundamental Identity (F.I.). However, the F.I. is typically violated under matrix regularization.
@@ 2.4 Restricted VPD (RVPD)
Physically, these constraints correspond to viewing the system in a uniformly accelerated frame. For a detailed discussion of this gauge restriction condition and its derivation, see our earlier paper[20].
@@ 2.6 Matrix-Regularized Supermembrane Action
These conditions ensure that the algebra closes consistently. For the detailed derivation of this algebraic structure and the proof of closure, we refer the reader to Appendix A and B of our previous work [20]. The condition $\partial_{3}\tau=0$ restricts the $\sigma^{3}$-dependence of the gauge transformation. However, this does not reduce the symmetry to purely static 2D diffeomorphisms; rather, it selects a specific subset of time-dependent transformations (akin to a uniformly accelerated frame [20]) that allows the Fundamental Identity to be satisfied in the matrix regularization.
@@ 4.1 Expansion of the Matrix Model
$S_{\mathrm{WZ}}=i\frac{T}{2}\int d\sigma^{3}\mathrm{Tr}\delta\bar{\theta}e^{\delta\bar{\theta}\delta_{S}}\left([X_{0}^{I},X_{0}^{J},\Gamma_{IJ}\delta\theta]+2[X_{0}^{I},\delta X^{J},\Gamma_{IJ}\delta\theta]\right).$ (4.6)
Here, the mixed triple commutators decompose as
$X_{0}^{I},X_{0}^{J};\delta X^{K}]=[\tau(X_{0}^{I},X_{0}^{J}),\delta X^{K}]+\frac{\partial\delta X^{K}}{\partial\sigma^{3}}[X_{0}^{I},X_{0}^{J}]+\Sigma(X_{0}^{I},X_{0}^{J};\delta X^{K}),$ (4.7)
$X_{0}^{I},\delta X^{J};\delta X^{K}]=[\tau(X_{0}^{I},\delta X^{J}),\delta X^{K}]+\frac{\partial\delta X^{K}}{\partial\sigma^{3}}[X_{0}^{I},\delta X^{J}]+\Sigma(X_{0}^{I},\delta X^{J};\delta X^{K}).$ (4.8)

### v3

@@ 2 Lorentz-Covariant M2-Brane Matrix Model with RVPD Gauge Symmetry
In this section, we review the Lorentz-covariant M2-brane matrix model with Restricted Volume-Preserving Deformations (RVPD), which we proposed in earlier work[https://doi.org/10.48550/arxiv.2504.05940, 20]. This review will serve as the basis for the BRST gauge-fixing procedure introduced in the next section.
@@ 2.2 Reformulation Using the Nambu Bracket
where $Q_{1},\ Q_{2}$ are arbitrary charges. The Poisson bracket property generalizes to the Nambu bracket, with the Jacobi identity replaced by the Fundamental Identity (F.I.). However, the F.I. is typically violated under matrix regularization. The Fundamental Identity is a defining property of the Nambu bracket in the continuum theory and ensures the consistency of the associated gauge symmetry [30]. It is well known that this identity is generically violated in finite-dimensional regularizations of the Nambu bracket, reflecting the truncation of the underlying infinite-dimensional algebra.
2.3 Algebraic preliminaries and notation
Throughout this paper we adopt the following notations:
@@ 2.4 Restricted VPD (RVPD)
Physically, these constraints correspond to viewing the system in a uniformly accelerated frame. For a detailed discussion of this gauge restriction condition and its derivation, see our earlier paper[https://doi.org/10.48550/arxiv.2504.05940].
@@ 2.6 Matrix-Regularized Supermembrane Action
These conditions ensure that the algebra closes consistently. For the detailed derivation of this algebraic structure and the proof of closure, we refer the reader to Appendix A and B of our previous work [18]. The condition $\partial_{3}\tau=0$ restricts the $\sigma^{3}$-dependence of the gauge transformation. However, this does not reduce the symmetry to purely static 2D diffeomorphisms; rather, it selects a specific subset of time-dependent transformations (akin to a uniformly accelerated frame [18] that allows the Fundamental Identity to be satisfied in the matrix regularization.
2.6.1 Equations of motion
Before discussing the BPS configurations, we summarize the equations of motion of the Lorentz-covariant RVPD matrix model.
The equations of motion are obtained by varying the matrix-regularized action with respect to the bosonic matrices $X^{I}$. In the bosonic sector, the Euler–Lagrange equation takes the form
$\epsilon^{ijk}\epsilon^{i^{\prime}j^{\prime}k^{\prime}}\left([\tau(X_{k^{\prime}},X_{j^{\prime}}),[X_{k},X_{j};X_{i}]]-2\,\partial_{\sigma_{3}}\big(X_{j^{\prime}}[X_{k^{\prime}},[X_{k},X_{i};X_{j}]]\big)\right)=0,$ (2.30)
which coincides with Eq. (5.2) of the purely bosonic Lorentz-covariant RVPD matrix model [18].
It should be emphasized that this equation is first order in the worldvolume time $\sigma_{3}$, reflecting the gauge-fixed membrane dynamics, but it represents the full second-order Euler-Lagrange equation derived from the action in the variational sense.
In the supersymmetric extension, the bosonic equation of motion is supplemented by the fermionic equation obtained from $\delta S/\delta\bar{\theta}=0$. The explicit form of the fermionic equation is omitted here, since it does not play an independent role at the classical level for the backgrounds considered. In this paper, we restrict ourselves to classical BPS backgrounds with $\theta_{0}=0$, so that the fermionic equation is trivially satisfied at the background level. Its linearization around the BPS background yields the Dirac-type operator appearing in the quadratic fermionic action used in the one-loop stability analysis.
The BPS solutions discussed in the next subsection satisfy first-order conditions originating from the $\kappa$-symmetry projection, and they automatically solve the above Euler-Lagrange equation.
@@ 2.7 Classical BPS Solutions
[…]
The noncommutative membrane configurations play the role of extended classical backgrounds in the Lorentz-covariant RVPD matrix model. They describe BPS membrane solutions that are two-dimensional objects embedded in eight transverse directions, providing physically well-defined vacua around which quantum fluctuations can be systematically analyzed. In this sense, these configurations represent membrane degrees of freedom encoded in the matrix model, rather than fundamental spacetime objects by themselves. While the underlying formulation is ten-dimensionally Lorentz covariant, the classical BPS membrane solutions extend only along these transverse directions. If a noncommutative membrane configuration is found to be unstable at the one-loop level, this indicates the presence of tachyonic fluctuation modes around the corresponding background. Such an instability is naturally interpreted as a dynamical rearrangement of matrix degrees of freedom toward a more stable BPS configuration, such as lower-dimensional membranes or particle-like states. A detailed description of the decay process lies beyond the scope of the present work and is left for future investigation.
@@ 4.1 Expansion of the Matrix Model
$S_{\mathrm{WZ}}=i\frac{T}{2}\int d\sigma^{3}\mathrm{Tr}\delta\bar{\theta}e^{\delta\bar{\theta}\delta_{S}}\left([X_{0}^{I},X_{0}^{J};\Gamma_{IJ}\delta\theta]+2[X_{0}^{I},\delta X^{J};\Gamma_{IJ}\delta\theta]\right).$ (4.6)
Using the decomposition of the triple bracket defined in Sec. 2.6, the mixed terms can be written as
$X^{I}_{0},X^{J}_{0};\delta X^{K}],\qquad[X^{I}_{0},\delta X^{J};\delta X^{K}].$ (4.7)